# Particle transport and deposition in wall-sheared thermal turbulence

### v1

@@ 3.1 Particle transport in the convection cell
We quantitatively describe the influence of wall shear on particle motion by analyzing the p.d.f.s of their velocity components ($u_{p}$ and $v_{p}$), as shown in figure 6. For particles with small and medium $St$ (see figures 6a,b), the p.d.f.s of $u_{p}$ are characterized by a single peak centered around zero, exhibiting symmetry for $Re_{w}\lesssim 5000$, suggesting a substantial likelihood of particles circulating within the LSC of convective flows. As $Re_{w}$ increases to 6000, the p.d.f.s of $u_{p}$ change to a bimodal distribution with two peaks off zero value, which can be attributed to the horizontal drift of particles caused by the wall shear. In addition, with increasing $St$, the peak corresponding to the negative $u_{p}$ exceeds that of the positive, suggesting an increased likelihood for particle suspension in the lower domain of the cell due to sedimentation (note the bottom wall is moving in the negative $x$-direction). For large $St$ (see figure 6c), we can deduce these particles are still minorly influenced by the turbulent flow because they exhibit horizontal motions with finite positive or negative $u_{p}$. Regarding the p.d.f.s of $v_{p}$, for small and medium $St$ (see figures 6d,e), their shapes are much narrower at $Re_{w}$ = 6000 compared to those at lower $Re_{w}$, indicating attenuated vertical motions in the zonal flow. This trend becomes more pronounced for large $St$ (see figure 6f), where $v_{p}$ values are predominantly negative. This implies vertical updrafts are insufficient to counter the gravitational settling, preventing these particles from rising with the flow currents.
@@
Figure 6: The p.d.f.s of particle velocity component (a-c) in the horizontal direction ($u_{p}$), and (d-f) in the vertical direction ($v_{p}$) with $Ra=10^{8}$
. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d) ${d_{p}}=5$ $\mu$m, (b,e) ${d_{p}}=20$ $\mu$m, and (c,f) ${d_{p}}=80$ $\mu$m.
@@
We examine the p.d.f.s of the particle Reynolds number $Re_{p}={d_{p}}|{{\boldsymbol{u}}_{f}}-{{\boldsymbol{u}}_{p}}|/{\nu_{f}}$, as shown in figure 7. We can see from figures 7($a,b$) that for particles with small $St$, the $Re_{p}$ is generally less than $10^{-3}$; for medium $St$, the $Re_{p}$ is generally less than $10^{-1}$. In the presence of a pronounced LSC within the flow, the p.d.f.s of the $Re_{p}$ manifest dual peaks, corresponding to the particles ascending or descending with the LSC. As wall shear increases, the $Re_{p}$ will also increase and exhibit enhanced intermittency. However, when $Re_{w}$ further increases up to 6000, the $Re_{p}$ decreases and their p.d.f.s change from a bimodal to unimodal distribution. This shift is attributed to the decreased vertical velocity component $v_{p}$ of the particles within the convective flow, as evidenced by figures 6(d-f). For large $St$, the $Re_{p}$ is as large as $Re_{p}\sim O(1)$ and spans a more extensive range, as shown in figure 7(c), which emphasizes the necessity of incorporating a drag correction factor $f(Re_{p})$ into the drag force calculations. We further plot the p.d.f.s of normalized particle Reynolds number $(Re_{p}-{\mu_{Re_{p}}})/{\sigma_{Re_{p}}}$, as shown in figures 7(d-f). For particles with small $St$ (see figure 7d), the p.d.f.s of $Re_{p}$ are close to the Gaussian distribution (denoted by the dashed line in the figure). However, with an increase in $Re_{w}$, the deviations in the tails of p.d.f.s. become more pronounced, indicating more intermittency events due to the influence of wall shear. For medium $St$ (see figure 7e), the p.d.f.s generally follow the Gaussian distribution, although there appear to be deviations with heavier tails, which implies there are more instances of $Re_{p}$ values deviating from the mean. For large $St$ (see figure 7f), the p.d.f.s of $Re_{p}$ are distinctly different from a Gaussian distribution, indicating a distinct distribution pattern where particle motion is nearly unaffected by the turbulence.
@@
Figure 7: (a-c) The p.d.f.s of particle Reynolds number $Re_{p}$, and (d-f) the p.d.f.s of normalized particle Reynolds number $(Re_{p}-{\mu_{Re_{p}}})/{\sigma_{Re_{p}}}$ with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d) ${d_{p}}=5$ $\mu$m, (b,e) ${d_{p}}=20$ $\mu$m, and (c,f) ${d_{p}}=80$ $\mu$m. Dashed lines in (d-f) represent the Gaussian distribution.
@@
Previous studies have shown that even in homogeneous isotropic turbulence, the particles may not distribute homogeneously but exhibit clustering behavior (Wang & Maxey, 1993; Bosse et al., 2006; Calzavarini et al., 2008). To identify the clustering behavior of particles during their transport in the wall-sheared convection cell, we employ a quantitative analysis using Voronoï diagrams (Monchaux et al., 2010, 2012). A Voronoï cell is defined as a collection of points closer to its corresponding particle than to any other; thus, the area (or volume) of a Voronoï cell inversely correlates with local particle concentration. In figure 8, we show Voronoï diagrams for the particle distribution, corresponding to the instantaneous state presented in figure 2. For particles with small $St$ (see figures 8a,d,g), Voronoï cells are of similar size and relatively evenly distributed in space, indicating a low degree of aggregation. For medium $St$ (see figures 8b,e,h), the areas of their Voronoï polygon exhibit much higher variability, indicating an increased level of spatial inhomogeneity. For large $St$ (see figures 8c,f,i), large voids are formed at the upper part of the convection cell, containing few to no particles. This is because most particles settle down and do not circulate in the cell, as revealed by the predominantly negative $v_{p}$ velocity components shown in figures 6(f).
@@
Figure 8: Typical instantaneous Voronoï diagrams for particles at (a-c) $Re_{w}=0$, (d-f) $Re_{w}=2000$, (g-i) $Re_{w}=6000$ with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m, and (c,f,i) ${d_{p}}=80$ $\mu$m.
@@
We then analyze the statistical distribution of the Voronoï cell area $A$. Figure 9 shows the p.d.f.s of the normalized Voronoï cell area $A/\left\langle A\right\rangle$, corresponding to the instantaneous state presented in figure 8. Here, $\left\langle A\right\rangle$ represents the mean value of $A$ at the current moment. For particles with small $St$ (see figure 9a), the Voronoï cells are randomly distributed throughout the flow field, and the distribution of their areas satisfies the $\Gamma$ distribution (Ferenc & Néda, 2007; Tagawa et al., 2012)
@@
where $d=2$ denotes the analysis in two-dimensions. For medium $St$ (see figure 9b), the p.d.f.s have a peak near unity, yet exhibit pronounced tails indicative of larger Voronoï cell areas, deviating from the $\Gamma$ distribution. This deviation from the $\Gamma$ distribution signals a departure from randomness, indicative of local preferential concentration. For large $St$ (see figure 9c), the p.d.f.s exhibit more pronounced tails towards larger Voronoï cell areas, indicating intensified void formation.
Figures 9(d-f) further show the time evolution of the standard deviation $\sigma$ of the normalized Voronoï cell area, serving as an indicator of spatial inhomogeneity. Due to particle deposition on the wall and constantly reducing particle numbers, the particles statistics cannot reach a statistically stationary state. Here, we only plot the time series from the initial state to the time at which half of the total particles remain suspended in the convection cell. For particles with small $St$ (see figure 9d), the standard deviation is the lowest, approaching the theoretical expectation for the $\Gamma$ distribution (i.e., $\sigma=0.53$, denoted by the dash-dotted line). For medium $St$ (see figure 9e), the greatest clustering is observed as the standard deviation of their Voronoï area significantly exceeds 0.53. For large $St$ (see figure 9f), the clustering also occurs; however, the influence of variations in wall shear strength $Re_{w}$ appears to be negligible in contrast to those with medium $St$.
Figure 9: (a-c) The p.d.f.s of the normalized Voronoï cell area $A/\langle A\rangle$ corresponding to the instantaneous state presented in figure 8, and (d-f) time evolution of the standard deviation of the normalized Voronoï cell area. (a,d) ${d_{p}}=5$ $\mu$m, (b,e) ${d_{p}}=20$ $\mu$m, and (c,f) ${d_{p}}=80$ $\mu$m.
Based on the above analysis, we arrive at the following mathematical model for the spatial distribution of particle concentration. We denote the initial concentration of particles as ${n_{0}}={N_{0}}/(LHW)$, where $N_{0}$ is the initial number of particles laden in the fluid. We define the instantaneous particle line density averaged over a horizontal cross-section is $n(\tau,y)=N(\tau,y)/(LW)$ at a given time $\tau$ and height $y$, where $N(\tau,y)$ is the instantaneous particle number at time $\tau$ and height $y$. We also define the time-averaged local particle concentration over a horizontal cross-section and along the cell height as $\langle n(y)\rangle_{t}$, where the average $\langle\cdot\rangle_{t}$ is calculated over time $t$. In transient processes, the particle concentration changes over time, and the time-averaged concentration represents the overall behavior of the particles during a specified time interval, which is crucial for understanding the long-term behavior of the system, even when the process is not in a steady state.
@@
For large $St$, sedimentation occurs almost independently of the carrier flow, analogous to their settling behavior in quiescent fluid environments. Thus, we can ignore their lateral motion and simply assume they settle at a constant speed of terminal velocity $v_{t}$. From the initial state of uniform distribution, after a duration of $\tau$, the particle ensemble settles a vertical distance of $v_{t}\tau$. Thus, the instantaneous particle line density $n(\tau,y)$ is
@@
We consider at the time $\tilde{t}\in\left[{0,t}\right]$, the instantaneous local particle concentration $n(\tilde{t},y)$ decreases from ${n_{0}}$ to 0 across a height interval $\tilde{y}$, where $\tilde{y}\in\left[{H-{v_{t}}\tilde{t},H}\right]$, resulting in $H-{v_{t}}\tilde{t}=\tilde{y}$. Thus, the horizontal- and time-averaged local particle concentration (further normalized by ${n_{0}}$), for large $St$, is
@@
In figure 10, we compare the theoretical estimations of ${\left\langle{n\left(y\right)}\right\rangle_{t_{1/2}}}/{n_{0}}$, namely (22) and (23), with both 2-D and 3-D simulation results. For particles with small $St$ (see figures 10a and 10d), the local particle concentration generally agrees with the theoretical predictions except for the discrepancy at $y>0.8H$. This discrepancy may be attributed to the finite $St$ effect (i.e., a small but not nonzero $St$), which results in a reduced particle presence in the upper boundary layer. At $Re_{w}=6000$ in the 2-D simulations, the flow transitions from the LSC to zonal flow, thus the assumption of spatially uniform particle concentration proposed by Martin & Nokes (1989) may no longer be valid, leading to significant deviation from (22). Specifically, when the LSC transitions to a zonal flow state, the updrafts are significantly weakened. This weakening results in fewer particles being drifted upwards and more particles settling down, leading to a higher particle concentration at the bottom of the zonal flow compared to the LSC state. For medium $St$ (see figures 10b and 10e), the absence of a corresponding model precludes a direct comparison; however, we provide a qualitative description of particle concentration. When the LSC is present, the local particle concentration appears to be consistent in the bulk region, but exhibits an increase with increasing $Re_{w}$ in the bottom boundary layer, implying that wall shear accelerates particle sedimentation. At $Re_{w}=6000$ in the 2-D simulations, the changes in flow states lead to distinctly different profiles for the local particle concentration compared to those at lower $Re_{w}$. For large $St$ (see figures 10c and 10f), the local particle concentration generally agrees with the theoretical prediction (23) across all $Re_{w}$, highlighting again that the particles motion is nearly not affected by the turbulent flows of the carrier fluid (i.e., either LSC or zonal flow state). When the wall shear becomes sufficiently strong, the vertical flow velocity component weakens significantly. Since the settling velocity of particles is strongly influenced by the vertical flow velocity component, the settling velocity of particles then approaches that in a stationary flow field (i.e., the terminal velocity $v_{t}$). This explains why the concentration profile for $Re_{w}=6000$ in figure 10(b) is similar to those in figure 10(c) for the 2-D simulation.
Figure 10: Vertical profiles of the horizontal- and time-averaged local particle concentration obtained from (a-c) the 2-D simulation with $Ra=10^{8}$ and (d-f) the 3-D simulation with $Ra=10^{7}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d) ${d_{p}}=5$ $\mu$m, (b,e) ${d_{p}}=20$ $\mu$m, and (c,f) ${d_{p}}=80$ $\mu$m. The pink solid line in (a,d) corresponds to equation (22), and the blue solid line in (c,f) corresponds to equation (23).

### v2

@@ 3.1 Particle transport in the convection cell
We quantitatively describe the influence of wall shear on particle motion by analysing the p.d.f.s of their velocity components ($u_{p}$ and $v_{p}$), as shown in figure 6. For particles with small and medium $St$ (see figures 6a,b), the p.d.f.s of $u_{p}$ are characterized by a single peak centred around zero, exhibiting symmetry for $Re_{w}\lesssim 5000$, suggesting a substantial likelihood of particles circulating within the LSC of convective flows. As $Re_{w}$ increases to 6000, the p.d.f.s of $u_{p}$ change to a bimodal distribution with two peaks off zero value, which can be attributed to the horizontal drift of particles caused by the wall shear. In addition, with increasing $St$, the peak corresponding to the negative $u_{p}$ exceeds that of the positive, suggesting an increased likelihood for particle suspension in the lower domain of the cell due to sedimentation (note the bottom wall is moving in the negative $x$-direction). For large $St$ (see figure 6c), we can deduce that these particles are still minorly influenced by the turbulent flow because they exhibit horizontal motions with finite positive or negative $u_{p}$. Regarding the p.d.f.s of $v_{p}$, for small and medium $St$ (see figures 6d,e), their shapes are much narrower at $Re_{w}$ = 6000 compared with those at lower $Re_{w}$, indicating attenuated vertical motions in the zonal flow. This trend becomes more pronounced for large $St$ (see figure 6f), where $v_{p}$ values are predominantly negative. This implies vertical updrafts are insufficient to counter the gravitational settling, preventing these particles from rising with the flow currents.
@@
Figure 6: The p.d.f.s of the particle velocity component (a-c) in the horizontal direction ($u_{p}$), and (d-f) in the vertical direction ($v_{p}$) with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d) ${d_{p}}=5$ $\mu$m, (b,e) ${d_{p}}=20$ $\mu$m and (c,f) ${d_{p}}=80$ $\mu$m.
@@
We examine the p.d.f.s of the particle Reynolds number $Re_{p}={d_{p}}|{{\boldsymbol{u}}_{f}}-{{\boldsymbol{u}}_{p}}|/{\nu_{f}}$, as shown in figure 7. We can see from figures 7($a,b$) that, for particles with small $St$, $Re_{p}$ is generally less than $10^{-3}$; for medium $St$, $Re_{p}$ is generally less than $10^{-1}$. In the presence of a pronounced LSC within the flow, the p.d.f.s of $Re_{p}$ manifest dual peaks, corresponding to the particles ascending or descending with the LSC. As wall shear increases, the $Re_{p}$ will also increase and exhibit enhanced intermittency. However, when $Re_{w}$ further increases up to 6000, $Re_{p}$ decreases and the p.d.f.s of $Re_{p}$ change from a bimodal to unimodal distribution. This shift is attributed to the decreased vertical velocity component $v_{p}$ of the particles within the convective flow, as evidenced by figures 6(d-f). For large $St$, $Re_{p}$ is as large as $Re_{p}\sim O(1)$ and spans a more extensive range, as shown in figure 7(c), which emphasizes the necessity of incorporating a drag correction factor $f(Re_{p})$ into the drag force calculations. We further plot the p.d.f.s of normalized particle Reynolds number $(Re_{p}-{\mu_{Re_{p}}})/{\sigma_{Re_{p}}}$, as shown in figures 7(d-f). For particles with small $St$ (see figure 7d), the p.d.f.s of $Re_{p}$ are close to the Gaussian distribution (denoted by the dashed line in the figure). However, with an increase in $Re_{w}$, the deviations in the tails of the p.d.f.s. become more pronounced, indicating more intermittency events due to the influence of wall shear. For medium $St$ (see figure 7e), the p.d.f.s generally follow the Gaussian distribution, although there appear to be deviations with heavier tails, which implies there are more instances of $Re_{p}$ values deviating from the mean. For large $St$ (see figure 7f), the p.d.f.s of $Re_{p}$ are distinctly different from a Gaussian distribution, indicating a distinct distribution pattern where particle motion is nearly unaffected by the turbulence.
@@
Figure 7: (a-c) The p.d.f.s of particle Reynolds number $Re_{p}$ and (d-f) the p.d.f.s of normalized particle Reynolds number $(Re_{p}-{\mu_{Re_{p}}})/{\sigma_{Re_{p}}}$ with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d) ${d_{p}}=5$ $\mu$m, (b,e) ${d_{p}}=20$ $\mu$m and (c,f) ${d_{p}}=80$ $\mu$m. Dashed lines in (d-f) represent the Gaussian distribution.
@@
Previous studies have shown that, even in homogeneous isotropic turbulence, the particles may not distribute homogeneously but exhibit clustering behaviour (Wang & Maxey, 1993; Bosse et al., 2006; Calzavarini et al., 2008). To identify the clustering behaviour of particles during their transport in the wall-sheared convection cell, we employ a quantitative analysis using Voronoï diagrams (Monchaux et al., 2010, 2012). A Voronoï cell is defined as a collection of points closer to its corresponding particle than to any other; thus, the area (or volume) of a Voronoï cell inversely correlates with local particle concentration. In figure 8, we show Voronoï diagrams for the particle distribution, corresponding to the instantaneous state presented in figure 2. For particles with small $St$ (see figures 8a,d,g), Voronoï cells are of similar size and relatively evenly distributed in space, indicating a low degree of aggregation. For medium $St$ (see figures 8b,e,h), the areas of their Voronoï polygon exhibit much higher variability, indicating an increased level of spatial inhomogeneity. For large $St$ (see figures 8c,f,i), large voids are formed at the upper part of the convection cell, containing few to no particles. This is because most particles settle down and do not circulate in the cell, as revealed by the predominantly negative $v_{p}$ velocity components shown in figures 6(f).
@@
Figure 8: Typical instantaneous Voronoï diagrams for particles at (a-c) $Re_{w}=0$, (d-f) $Re_{w}=2000$, (g-i) $Re_{w}=6000$ with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m and (c,f,i) ${d_{p}}=80$ $\mu$m.
@@
We then analyse the statistical distribution of the Voronoï cell area $A$. Figure 9 shows the p.d.f.s of the normalized Voronoï cell area $A/\left\langle A\right\rangle$, corresponding to the instantaneous state presented in figure 8. Here, $\left\langle A\right\rangle$ represents the mean value of $A$ at the current moment. For particles with small $St$ (see figure 9a), the Voronoï cells are randomly distributed throughout the flow field, and the distribution of their areas satisfies the $\Gamma$ distribution (Ferenc & Néda, 2007; Tagawa et al., 2012)
@@
where $d=2$ denotes the analysis in two dimensions. For medium $St$ (see figure 9b), the p.d.f.s have a peak near unity, yet exhibit pronounced tails indicative of larger Voronoï cell areas, deviating from the $\Gamma$ distribution. This deviation from the $\Gamma$ distribution signals a departure from randomness, indicative of local preferential concentration. For large $St$ (see figure 9c), the p.d.f.s exhibit more pronounced tails towards larger Voronoï cell areas, indicating intensified void formation.
Figures 9(d-f) further shows the time evolution of the standard deviation $\sigma$ of the normalized Voronoï cell area, serving as an indicator of spatial inhomogeneity. Due to particle deposition on the wall and constantly reducing particle numbers, the particles statistics cannot reach a statistically stationary state. Here, we only plot the time series from the initial state to the time at which half of the total particles remain suspended in the convection cell. For particles with small $St$ (see figure 9d), the standard deviation is the lowest, approaching the theoretical expectation for the $\Gamma$ distribution (i.e. $\sigma=0.53$, denoted by the dash-dotted line). For medium $St$ (see figure 9e), the greatest clustering is observed as the standard deviation of their Voronoï area significantly exceeds 0.53. For large $St$ (see figure 9f), the clustering also occurs; however, the influence of variations in wall-shear strength $Re_{w}$ appears to be negligible in contrast to those with medium $St$.
Figure 9: (a-c) The p.d.f.s of the normalized Voronoï cell area $A/\langle A\rangle$ corresponding to the instantaneous state presented in figure 8, and (d-f) time evolution of the standard deviation of the normalized Voronoï cell area. Panels show (a,d) ${d_{p}}=5$ $\mu$m, (b,e) ${d_{p}}=20$ $\mu$m and (c,f) ${d_{p}}=80$ $\mu$m.
Based on the above analysis, we arrive at the following mathematical model for the spatial distribution of particle concentration. We denote the initial concentration of particles as ${n_{0}}={N_{0}}/(LHW)$, where $N_{0}$ is the initial number of particles laden in the fluid. We define the instantaneous particle line density averaged over a horizontal cross-section as $n(\tau,y)=N(\tau,y)/(LW)$ at a given time $\tau$ and height $y$, where $N(\tau,y)$ is the instantaneous particle number at time $\tau$ and height $y$. We also define the time-averaged local particle concentration over a horizontal cross-section and along the cell height as $\langle n(y)\rangle_{t}$, where the average $\langle\cdot\rangle_{t}$ is calculated over time $t$. In transient processes, the particle concentration changes over time, and the time-averaged concentration represents the overall behaviour of the particles during a specified time interval, which is crucial for understanding the long-term behaviour of the system, even when the process is not in a steady state.
@@
For large $St$, sedimentation occurs almost independently of the carrier flow, analogous to their settling behaviour in quiescent fluid environments. Thus, we can ignore their lateral motion and simply assume they settle at a constant speed of terminal velocity $v_{t}$. From the initial state of uniform distribution, after a duration of $\tau$, the particle ensemble settles a vertical distance of $v_{t}\tau$. Thus, the instantaneous particle line density $n(\tau,y)$ is
@@
We consider at the time $\tilde{t}\in\left[{0,t}\right]$ the instantaneous local particle concentration $n(\tilde{t},y)$ that decreases from ${n_{0}}$ to 0 across a height interval $\tilde{y}$, where $\tilde{y}\in\left[{H-{v_{t}}\tilde{t},H}\right]$, resulting in $H-{v_{t}}\tilde{t}=\tilde{y}$. Thus, the horizontal- and time-averaged local particle concentration (further normalized by ${n_{0}}$), for large $St$, is
@@
In figure 10, we compare the theoretical estimations of ${\left\langle{n\left(y\right)}\right\rangle_{t_{1/2}}}/{n_{0}}$, namely (22) and (23), with both 2-D and 3-D simulation results. For particles with small $St$ (see figures 10a,d), the local particle concentration generally agrees with the theoretical predictions except for the discrepancy at $y>0.8H$. This discrepancy may be attributed to the finite $St$ effect (i.e. a small but not non-zero $St$), which results in a reduced particle presence in the upper boundary layer. At $Re_{w}=6000$ in the 2-D simulations, the flow transitions from the LSC to zonal flow, thus the assumption of a spatially uniform particle concentration proposed by Martin & Nokes (1989) may no longer be valid, leading to significant deviation from (22). Specifically, when the LSC transitions to a zonal flow state, the updrafts are significantly weakened. This weakening results in fewer particles drifting upwards and more particles settling down, leading to a higher particle concentration at the bottom of the zonal flow compared with the LSC state. For medium $St$ (see figures 10b,e), the absence of a corresponding model precludes a direct comparison; however, we provide a qualitative description of particle concentration. When the LSC is present, the local particle concentration appears to be consistent in the bulk region, but exhibits an increase with increasing $Re_{w}$ in the bottom boundary layer, implying that wall shear accelerates particle sedimentation. At $Re_{w}=6000$ in the 2-D simulations, the changes in flow states lead to distinctly different profiles for the local particle concentration compared with those at lower $Re_{w}$. For large $St$ (see figures 10c,f), the local particle concentration generally agrees with the theoretical prediction (23) across all $Re_{w}$, highlighting again that the particles’ motion is nearly not affected by the turbulent flows of the carrier fluid (i.e. either LSC or zonal flow state). When the wall shear becomes sufficiently strong, the vertical flow velocity component weakens significantly. Since the settling velocity of particles is strongly influenced by the vertical flow velocity component, the settling velocity of particles then approaches that in a stationary flow field (i.e. the terminal velocity $v_{t}$). This explains why the concentration profile for $Re_{w}=6000$ in figure 10(b) is similar to those in figure 10(c) for the 2-D simulation.
Figure 10: Vertical profiles of the horizontal- and time-averaged local particle concentration obtained from (a-c) the 2-D simulation with $Ra=10^{8}$ and (d-f) the 3-D simulation with $Ra=10^{7}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d) ${d_{p}}=5\ \mu$m, (b,e) ${d_{p}}=20\ \mu$m and (c,f) ${d_{p}}=80\ \mu$m. The pink solid line in (a,d) corresponds to (22), and the blue solid line in (c,f) corresponds to (23).